OpenStudy (anonymous):

the sum of two numbers is seven. the sum of the squares of the two numbers is twenty nine. find the two numbers

OpenStudy (anonymous):

The two numbers are x and y x + y = 7 (1st sentence) x^2 + y^2 = 29 (2nd sentence) two equations two variables... You could solve by graphing, substitution, or elimination method. Which way are you solving them right now?

OpenStudy (anonymous):

substitution

OpenStudy (anonymous):

so 25^2+4^2=29 ?

OpenStudy (anonymous):

opps i meant 5 not 25

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

x + y = 7 when I solve for x I get x = 7 - y substitute that into the second equation (7 - y)^2 + y^2 = 29 foil method the first binomial (7 - y)(7 - y) + y^2 = 29 49 - 7y - 7y + y^2 + y^2 = 29 Add like terms 2y^2 - 14y + 49 = 29 set it equal to 0 2y^2 - 14y + 20 = 0 factor (take out the 2 first) 2(y^2 - 7y + 10) = 0 factor 2(y - 5)(y - 2) = 0 set each factor = to 0 y - 5 = 0 or y - 2 = 0 solve each equation y = 5 or y = 2 if y = 5 then x = 2 because x + y = 7 if x = 5 then y = 2 because x + y = 7 so the two numbers are 2 and 5

OpenStudy (anonymous):

yes thanks again